OpenStudy (anonymous):

Point V is graphed on a number line at –6. Point U is 8 units away from point V on the number line. What could be the coordinates of point U? Choose all answers that are correct. A. –14 B. –2 C. 1 D. 2

OpenStudy (anonymous):

help

OpenStudy (igreen):

Add -6 + 8 and subtract -6 - 8, what two numbers do you get?

OpenStudy (igreen):

Can you do that? @newwar248

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Okay, what do you get?

OpenStudy (igreen):

-6 + 8 = ? -6 - 8 = ?

OpenStudy (anonymous):

14 and 2

OpenStudy (igreen):

Not quite. -6 - 8 = -14..not 14. So -14 and 2 are your answers :)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @newwar248

OpenStudy (anonymous):

do you mind one more

OpenStudy (anonymous):

Which equation demonstrates the opposite of a sum property? A. (22 + 10) + 3 = (10 + 22) + 3 B. –5 + 13 = 13 + (–5) C. (12 + (–4)) + 14 = 12 + ((–4) + 14) D. –(–6 + 17) = 6 + (–17)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

I'm not sure, but I think it's B.

OpenStudy (anonymous):

thanks